OpenStudy (fly_g):

Two high divers jump off different platforms at the same time. The height of the first diver, in feet, is given by f(x) = –16x2 + 9x + 67, while the height of the second diver is given by f(x) = –16x2 + 15x + 54, where x is the number of seconds after they jump. When are the divers at the same height? Round to the nearest hundredth of a second.

OpenStudy (kamibug):

You can set the two equations equal to each other b/c they represent the height of each diver and you're looking for when they're at the same height. :) -16x^2 + 9x + 67 = -16x^2 + 15x + 54 When you add -16x^2 to both sides it cancels, so now you have: 9x + 67 = 15x + 54 Can you solve for x from here? :) \(\Huge\bf\color{seagreen}{Welcome~to~Openstudy!~:)}\)

OpenStudy (fly_g):

yes thank you, ill try.

OpenStudy (fly_g):

wait do i cancel out the x's first or the regular numbers?

OpenStudy (kamibug):

You want to isolate x. Combine Like Terms by subtracting 9x from both sides and subtracting 54 from both sides. 9x + 67 = 15x + 54 13 = 6x Now divide both sides by 6 to solve for x. x = approx. 2.17 seconds

OpenStudy (fly_g):

thank you, so if i was rounding to the nearest hundreth the answer would be 2.20?

OpenStudy (fly_g):

or 2.17

OpenStudy (kamibug):

Well, I already rounded, Lol. It was 2.16 repeating, so it rounds to 2.17. :)